# Graph RAG for Legal Norms: A Hierarchical and Temporal Approach

### v2

@@ 3.6 Hierarchical Grouping as Communities
The grouping components of legal norms function in practice as hierarchical levels for thematic grouping, chosen as the systematizing criterion for the articulation. As observed in the example of Figure 4, Art. 5º (Article 5) is grouped under “CAPÍTULO I – DOS DIREITOS E DEVERES INDIVIDUAIS E COLETIVOS” (CHAPTER I – OF INDIVIDUAL AND COLLECTIVE RIGHTS AND DUTIES), which in turn is grouped under “TÍTULO II - DOS DIREITOS E GARANTIAS FUNDAMENTAIS” (TITLE II - OF FUNDAMENTAL RIGHTS AND GUARANTEES). These groupings then ascend in level until reaching the Norm (as a Whole), representing natural hierarchical and thematic structures within legal norms. This directly aligns with the concept of communities in a graph proposed by the Graph RAG approach – in this approach, “communities” refer to groups of nodes in the graph that are more densely interconnected among themselves than with the rest of the graph, representing thematic or conceptual clusters. Therefore, the hierarchical structure of legal norms can be viewed as an intrinsic form of organization into thematic communities.
@@
Legal norms, in general, are classified into one or more themes [^21] Brazilian federal norms are classified according to the “Sistema de Classificação de Normas Jurídicas do Congresso Nacional” (System for Classification of Legal Norms of the National Congress), with examples of themes including: “Children and Adolescents”, “Extraordinary Credit”, “Indigenous Population”, “Basic Sanitation”, ’General Social Security System’, etc., which can be considered intertextual, or inter-norm, aggregators, transversally gathering norms on cohesive subjects, as exemplified in Figure 17. In this figure, the Norm (as a whole) are represented by circles in green, and the Themes, in orange.
@@
Figure 17: Diagram illustrating the inter-norm aggregation by legal themes (orange), represented as higher-level communities that group "Norms" (green) in the knowledge graph.
Components of legal norms could also be classified under one or more themes, as demonstrated in Figure 18. This allows broadening the classification of the components beyond their groupers within the hierarchical structure of the legal norm.
@@
Figure 18: Diagram illustrating associations between components and legal themes.
The themes would then be inserted as Entities/Nodes Theme in the Knowledge Graph, each having an associated Text Unit with its description, which would be produced by a human domain expert, ensuring the curation of the produced texts.
Another approach could be to use an LLM to generate the theme descriptions, producing summaries from the descriptions of the sub-themes or, in the case of leaf themes, from the summaries of the associated norms - it would require intensive processing and high costs (considering that a theme can cover hundreds of norms). The LLM would act as an assistant in a semi-automatic generation, where these descriptions would be reviewed and potentially updated by the human domain expert - the summary’s effectiveness would heavily depend on the quality of the algorithms and models used.
3.7 Retrieval considering Structure
Another advantage of using a graph to represent structural entities, reflecting the hierarchical structure of legal texts, is that the user could choose the scope of their query (level of aggregation), for example, by defining that they wish to query about a specific theme (which would encompass all its sub-themes and associated norms, with their components and versions), or about a specific title or article of a norm (which would encompass all its subdivisions and versions), or about a specific version of a norm or a component (which would encompass its child versions), as presented in Figure 19.
Figure 19: Diagram illustrating the possibility of retrieving Text Units considering the hierarchical structure of the knowledge graph
The choice defines a filter on the hierarchical index of the Text Units (and vectors/embeddings) to be retrieved, making the context retrieval for the response more focused. The hierarchical structure of the graph allows for more precise semantic navigation between different levels of detail. This “semantic navigation” manifests in the user’s ability to direct the search to a specific level of the legal hierarchy (theme, title, chapter, article, etc.), filtering the search space and focusing retrieval on the Text Units most relevant to their interest at that level of detail. Instead of a generic search across the entire graph, the user can explore legal knowledge in a more structured and hierarchy-guided manner.
@@ 4 Conclusion and Future Work
Aiming to enhance the analysis and comprehension of legal norms in the era of Artificial Intelligence, this article presented a detailed proposal for the construction of a legal knowledge graph, enriched with contextually relevant Text Units. The central objective was to expand the scope of queries regarding legal norms, allowing Artificial Intelligence systems to address questions ranging from literal text retrieval to complex inquiries requiring an understanding of the hierarchical structure, the intrinsic semantics, the intra- and inter-relations and the temporal evolution of law.
Incorporating such structural and temporal knowledge is expected to enhance the precision of legal AI responses, thereby reducing contextual errors and the risk of citing incorrect versions of the law.
@@
  • A detailed proposal for building a legal knowledge graph, integrating the hierarchical structure, temporal versioning, and cumulative and contextually enriched Text Units, with an initial focus on structural entities.
  • The modeling of Versions as Aggregations, reflecting hierarchical interdependence and the propagation of changes over time, allowing for an accurate and efficient representation of normative evolution.
  • The suggestion to incorporate Additional Context into versions, enriching the graph with descriptive metadata about the history and lifecycle of norms, providing additional context during response generation by the LLM.
  • The proposal to generate specific Text Units for Alteration Actions (e.g., repeal, text amendment), making the norm-modifying events themselves and their effects semantically retrievable.
  • The generation of Text Units from structured Metadata (intrinsic properties and relationships such as succession or correlation), allowing non-explicitly textual information in norms to be included in vector search and aligning with concepts like multi-aspect embeddings.
  • The perspective of Hierarchical Grouping as Communities, interpreting the hierarchical structure and legal themes (curated by specialists) as thematic communities within the graph, contributing to more targeted and semantic information retrieval strategies.
  • The proposal of Retrieval considering Structure, enabling users to define the search scope and semantically navigate the legal hierarchy (whether by theme, title, article, version, etc.), refining the precision and relevance of responses.
  • A Graph RAG approach that seeks the lowest possible cost for graph creation by prioritizing pre-existing or curated hierarchical and thematic structures, avoiding computationally intensive techniques such as automatic community detection and the construction of complex summaries by LLMs for this purpose.
The main innovation lies in the adaptation and expansion of Graph RAG to the legal field, with a particular focus on structural entities, considered essential for identifying robust legal grounds. Semantic segmentation guided by the hierarchical structure, combined with the representation of cross-references, versioning and grouping into thematic communities, provides a natural and rich contextual basis for legal RAG strategies.
The proposal to use the hierarchy of communities and themes designed by domain experts reflects the original intent of legislators and ensures informational curation, an essential aspect for the reliability of AI systems in Law, in addition to minimizing the need for complex and costly graph construction processes.

### v3

@@ 3.5 Hierarchical Grouping as Communities
The hierarchical structure of a legal norm provides a natural, intrinsic form of community organization. The grouping elements described earlier—such as Titles, Chapters, and Sections—are not merely textual containers; they are abstract Component entities that provide thematic context. As observed in Figure 4, the Component "Art. 5º" is part of the "CAPÍTULO I" Component, which in turn is part of the "TÍTULO II" Component. This creates a nested hierarchy of Works that ascends to the main Norm.
This structure directly aligns with the concept of "communities" in Graph RAG, where communities are densely interconnected clusters of nodes. In our model, the predefined legal hierarchy is a primary, curated form of community structure, superior to purely algorithmic community detection.
@@
Beyond this internal hierarchy, legal norms are often classified into external, cross-cutting themes [^20] Brazilian federal norms are classified according to the “Sistema de Classificação de Normas Jurídicas do Congresso Nacional” (System for Classification of Legal Norms of the National Congress), with examples of themes including: “Children and Adolescents”, “Extraordinary Credit”, “Indigenous Population”, “Basic Sanitation”, ’General Social Security System’, etc.. These themes can be modeled as high-level community Theme entities (nodes) that group multiple Norm entities (Works) on cohesive subjects, creating inter-norm communities, as exemplified in Figure 17.
@@
Figure 17: Diagram illustrating inter-norm aggregation by legal Theme entities/nodes (orange), represented as higher-level communities that group Norm (Work) entities (green) in the knowledge graph.
This thematic classification can also be applied at a more granular level, linking Theme nodes directly to specific Component entities, as shown in Figure 18. This allows for a richer, multi-faceted classification of legal provisions beyond their formal position in the document hierarchy.
@@
Figure 18: Diagram illustrating associations between Component (Work) entities and legal Theme (Community).
These Theme nodes will be inserted into the Knowledge Graph, each with an associated Text Unit describing the theme’s scope (there can be a Text Unit for each language). This description could be produced by a human domain expert or generated semi-automatically with LLM assistance, where the LLM produces a summary based on the definitions of sub-themes or the content of the associated norms. The latter approach would require careful specialized human review, as the quality of the summary is critical. In either case, using predefined thematic structures provided by official sources is preferable, as it ensures the semantic grouping reflects established legal practice.
3.6 Structure-Aware Retrieval
A key advantage of our graph-based model is enabling structure-aware retrieval. By representing the full hierarchy of Themes, Norms, Components, and their Versions, the user can precisely define the scope of their query. For example, a user could choose to search within:
  • A specific Theme, retrieving Text Units from all sub themes, norms and components classified under it.
  • A specific Norm (e.g., the entire Civil Code), searching across all its components.
  • A specific Component (e.g., a single Title or Article), restricting the search to its subdivisions.
  • A specific Temporal Version of a norm or component, querying only the text as it existed on a particular date.
This process is illustrated in Figure 19.
Figure 19: Diagram illustrating how a user can select a scope (e.g., a Theme, Norm, Component or Version) to filter the retrieval of relevant Text Units from the knowledge graph.
Choosing a scope acts as a powerful filter on the set of Text Units to be considered for retrieval. Instead of performing a vector search across the entire legal corpus, the system first traverses the graph from the selected entry point (e.g., a "Title" Component) to identify your own Text Units and those of all your descendants. Only the Text Units belonging to this filtered subset are then used in the semantic search.
This hierarchical filtering allows for more precise "semantic navigation". The user can move between different levels of detail, from a broad thematic query to a focused analysis of a single article’s evolution. This structured, hierarchy-guided exploration of legal knowledge is more efficient and yields more relevant results than a flat search across an entire, undifferentiated database of text chunks.
@@ 4 Conclusion and Future Work
Aiming to enhance the analysis and comprehension of legal norms in the era of Artificial Intelligence, this article presented a detailed proposal for the construction of a legal knowledge graph. By grounding the graph’s design in a formal, FRBRoo-based model for temporal and linguistic versioning, we have outlined a robust methodology for representing the complex lifecycle of legal texts. The central objective was to expand the scope of queries regarding legal norms, allowing AI systems to address questions ranging from literal text retrieval to complex inquiries requiring a deterministic understanding of the hierarchical structure, intrinsic semantics, and temporal evolution of law.
@@
  • A formal, multi-layered model for legal knowledge graphs, distinguishing between abstract Works (Norms and Components) and their concrete textual Expressions (Temporal and Language Versions). This provides a solid theoretical foundation for representing legal documents.
  • The modeling of Temporal Versions as Aggregations, reflecting hierarchical interdependence and enabling an accurate and efficient representation of normative evolution by reusing unchanged components over time.
  • The proposal to generate specific Text Units for normative Events (Actions), making the norm-modifying events themselves and their effects (e.g., a change in validity) semantically retrievable.
  • The generation of Text Units from structured Metadata associated with abstract Norm and Component entities, allowing non-explicitly textual information to be included in vector search and aligning with multi-aspect embedding concepts.
  • The interpretation of the intrinsic legal hierarchy and extrinsic thematic classifications as curated Communities within the graph, enabling more targeted retrieval strategies.
  • The proposal of Structure-Aware Retrieval, which leverages the graph’s hierarchy to allow users to precisely define a search scope, refining the precision and relevance of responses.
  • A Graph RAG approach that seeks to minimize creation costs by prioritizing pre-existing, curated structures (both hierarchical and thematic) over computationally intensive, purely algorithmic methods.
The main innovation lies in the detailed adaptation and expansion of the Graph RAG paradigm for the legal field. By integrating a formal versioning model inspired by library science (FRBRoo), we provide a deterministic "ground truth" that is essential for reliable AI in a high-stakes domain like Law. This structured approach, combining a formal hierarchy, explicit versioning, and curated thematic communities, provides a rich and natural contextual basis for legal RAG strategies.
The proposal to use the hierarchy of communities and themes designed or reviewed by domain experts reflects the original intent of legislators and ensures informational curation, an essential aspect for the reliability of AI systems in Law, in addition to minimizing the need for complex and costly graph construction processes.